### v1

# Monte Carlo Simulation Development and Implementation of the GiBUU Model for Neutrino Experiments

Abstract
This paper introduces a Monte Carlo simulation generated with the GiBUU model for neutrino experiments. The simulation generates realistic neutrino event samples, contributing to the prediction and interpretation of experimental outcomes. The results showcase the performance of the GiBUU-based simulation framework, emphasizing its fidelity to the original GiBUU cross-section model. This first implementation enables future work on developing the infrastructure to propagate systematic uncertainties. These contributions enhance the precision of experimental predictions and provide a platform for further exploration in future studies.
preprint: APS/123-QED
I Introduction
Neutrinos, elusive and ghostly particles, have captivated the attention of physicists for decades. Neutrinos possess intriguing properties, such as their weak interactions and tiny masses. These characteristics make neutrinos fascinating subjects of study, with profound implications for our understanding of the fundamental laws of nature. Neutrino physics is essential in particle physics, offering unique insights into the fundamental forces and particles that govern the universe.
Neutrino masses are demonstrated because they oscillate between flavors (electron, muon, and tau) as they propagate through space, and this phenomenon provides unequivocal evidence for physics beyond the Standard Model [1, 2]. Neutrino oscillation experiments have led to groundbreaking discoveries, challenging our understanding of neutrino properties and paving the way for further investigations beyond the known realm.
In particular, accelerator-based neutrino experiments have significantly advanced our understanding of neutrinos and their properties. These experiments employ powerful particle accelerators to produce intense and well-characterized neutrino beams, allowing for detailed investigations of neutrino interactions, oscillations, and their potential implications for other physics beyond the Standard Model. Understanding neutrino oscillations is paramount for advancing our knowledge of the lepton sector and exploring new physics; for example, deviations from the observed oscillation patterns could indicate the presence of new neutrino flavors or interactions that go beyond the established framework of the Standard Model. Moreover, these experiments provide a unique avenue for probing the Universe’s asymmetry between matter and antimatter.
In this context, Monte Carlo simulations are essential for designing, analyzing, and interpreting experimental data. These simulations provide a means to model and predict neutrino interactions with target materials, detector responses, and backgrounds, thus enabling precise estimates of the expected event rates to extract the desired physics parameters, including corrections due to limited efficiencies and backgrounds. The importance of accurate models in Monte Carlo simulations stems from their significant impact on result extraction and the subsequent interpretation of physics outcomes.
This paper presents a Monte Carlo event generator based on the GiBUU (Giessen Boltzmann-Uehling-Uhlenbeck) model [3, 4] and specifically tailored for neutrino experiments. Within this work, the GiBUU model, renowned for its success in describing hadron/electron/photon/neutrino-nucleus interactions in various target materials, is converted into an event-by-event predictor. By incorporating the GiBUU model into a Monte Carlo simulation framework, we aim to provide a comprehensive tool for simulating and analyzing neutrino data events.
This implementation has been proved into the Neutrino Event Generation stage of LArSoft (Liquid Argon Software [5]) $v09\_53\_02$, and it has been tested using the SBND [6] and ICARUS [7] detector geometries. Both geometries are traversed by an enhanced $\nu_{\mu}$ beam from the BNB (Booster Neutrino Beam) beamline at Fermilab [8]. The BNB flux peaks at 0.6-0.8 GeV in these experiments. The distributions presented in this report were generated using a simulated box geometry with dimensions of 2 m on each side and filled with Liquid Argon, located 110 m from the BNB target, referred to as LArBox throughout the paper.
The simulation leverages the GENIE [9, 10] geometry and flux drivers and produces the output event formats compatible with those from GENIE. By adopting this approach, we can run experiment-specific software to reconstruct the generated events in the same manner as we would with GENIE-generated events, eliminating the burden of developing additional software at each experiment to read GiBUU Monte Carlo outputs. The relevance of our Monte Carlo generator extends beyond the accurate modeling of neutrino-nucleus cross sections; it enables precise predictions of neutrino event rates, energy spectra, and particle kinematic distributions. These predictions are crucial for the interpretation of oscillation measurements and the search for new physics phenomena. The model accounts for various aspects, including final-state interactions and nuclear effects, and the generated events can quickly propagate to simulate the detector response using experiment-specific software, reproducing observed experimental signatures accurately using the same software setup as neutrinos generated by GENIE Monte Carlo.
Section II offers a brief description of the GiBUU model that is relevant for neutrino-nucleus scattering in the context of neutrino oscillations. Section III provides an overview and discussion of how the Monte Carlo event generator is produced. A brief discussion of the following steps towards calculating and implementing the systematic uncertainties associated with the GiBUU model is presented in Section IV, and the results of the work and final remarks are presented in Section V.
II The GiBUU model and its relevance to neutrino interactions
The GiBUU (Giessen Boltzmann-Uehling-Uhlenbeck) [3, 4] model serves as a comprehensive theoretical framework designed to study interactions of hadrons, electrons, and neutrinos within various nuclear systems. Grounded in kinetic theory and effective quantum field theory principles, GiBUU incorporates nuclear transport theory and mean-field dynamics concepts to describe the evolution of hadronic systems by considering their interactions through scattering, decay, and production processes. The GiBUU model has found extensive use in a range of research areas, including nuclear reactions, heavy-ion collisions, and neutrino-nucleus interactions, the latter being the focus of this work. Its versatility and accuracy have made it valuable in interpreting experimental data and understanding the underlying dynamics of nuclear systems.
A essential feature of the GiBUU model is its remarkable accuracy in replicating various experimental datasets without the need for tuning to neutrino scattering data, in which several assumptions of the initial neutrino would have been to be made. To develop some of the model parametrizations, the GiBUU model contains tunings to electron and hadron experimental scattering data, where precise knowledge of the incoming particle energy is available. GiBUU has been extensively tested and validated against various measurements, demonstrating its stability and robustness in describing neutrino-nucleus interactions, including argon targets [11]. This unique feature of the GiBUU model highlights its inherent ability to capture the intricacies of neutrino interactions without relying on specific adjustments or tuning procedures to match neutrino experimental results. Therefore, using the GiBUU model for calculating the efficiency, the background description, and other model-dependent measurement corrections is expected to reduce associated systematic uncertainties. Furthermore, the advantage of not relying on neutrino data tuning allows for more robust handling of the related uncertainties, as comparisons of the model with neutrino scattering data can help estimate some of the model’s uncertainties. By leveraging the inherent accuracy and stability of the GiBUU model, the systematic uncertainties associated with neutrino experiments can be effectively addressed and minimized, enhancing the reliability and precision of the results obtained.
GiBUU treats the neutrino-nucleus interaction using weak charged current (CC) and neutral current (NC) interaction processes. It considers the quasi-elastic scattering (QE), resonant and non-resonant production mechanisms, deep inelastic scattering, and meson-exchange current processes that play significant roles in neutrino interactions.
The GiBUU model intricately incorporates nuclear effects via a comprehensive treatment of Final State Interactions (FSI) with nucleon-level interactions. FSI accounts for the intricate interplay between the produced hadrons and the encompassing nuclear medium, including rescattering absorption and re-emission processes. To meticulously capture these interactions with nuclei, the model integrates nuclear structure information, maintains a coherent nuclear potential description, and employs the transport model (the BUU equation). This approach ensures a precise treatment of nuclear effects and elucidates the nuanced influence of a nucleus on these interactions. In GiBUU, this facet proves indispensable for the comprehension of hadronic observables arising from the presence of nuclear matter. This aspect is crucial for neutrino experiments analyzing exclusive neutrino scattering topologies, in which the description of each of the particles produced leaving the nucleus plays an essential role in neutrino energy reconstruction.
The GiBUU model undergoes rigorous validation against diverse neutrino experimental datasets, encompassing charged-current and neutral-current interactions [12, 13]. Though tensions with some datasets have been noted—through prior data corrections using alternative Monte Carlo methods—GiBUU’s overarching performance remains notably robust. Its capability to faithfully replicate experimental observables, spanning cross-sections, energy spectra, and angular distributions, attests to its reliability and accuracy in portraying the intricate dynamics inherent to neutrino interactions. While recognizing localized disparities, the overarching proficiency of GiBUU positions it as a formidable tool for comprehensive analyses in neutrino physics.
III Monte Carlo Simulation Framework for Neutrino Experiments
One of the most common neutrino Monte Carlo event generators, GENIE [9, 10], has emerged as an essential tool for simulating neutrino interactions in various experiments, and in particular for liquid argon experiments at Fermilab, such as MicroBooNE [14], ICARUS [7], SBND [6] and DUNE [15] on which coordinated efforts for development, tuning, and uncertainty estimation of the GENIE generator are already formed. By generating simulated events that mimic real neutrino interactions, GENIE facilitates interpreting data obtained from neutrino experiments and aids in developing and optimizing experimental designs. The widespread utilization of the GENIE event generator in numerous neutrino experiments establishes its critical role. Consequently, it is unsurprising that any subsequent software infrastructure development encompassing detector simulation and event reconstruction is inherently compatible with the output format of GENIE for many experiments and, in particular, for USA-based experiments. Hence, within this work, the simulation of neutrino events using the GIBUU model, considering a neutrino beam and detector geometry (LArBox), utilizes the GENIE flux and geometry drivers. Furthermore, the resulting generated GiBUU events are formatted in accordance with GENIE’s format to ensure seamless integration and compatibility.
Our simulation consists of two distinct steps. Firstly, a significant number of GiBUU events are simulated utilizing the original GiBUU model framework that the Giessen group developed. We produced $1.3\times 10^{10}$ events aiming to generate extensive libraries of neutrino events. This approach ensures event uniqueness and minimizes repetitions for each neutrino flavor and interaction type based on a given neutrino flux. The original GiBUU simulation provides cross-sections, including interaction amplitudes, some carrying negative weights, signifying their contribution as destructive interference terms in the overall cross-section calculation. These negative weights arise from non-resonant pion background processes, particularly relevant at higher energies such as those expected at the DUNE experiment. Each entry of the libraries stores the kinematic information of the incoming neutrino and produced particle kinematics, its GiBUU weight, and the event generator reaction code, which denotes the initial neutrino-nucleus interaction mode. Two libraries are created; one contains events with positive weights and the other with negative weights. In the latter case, the negative weights are converted to positive by taking their absolute values.
Secondly, two independent productions are made, one per library set, using an acceptance-rejection method, event by event, to select the input from the library. The weight distribution from narrow neutrino energy in the library is used as probability distribution function input to the method, like the sample shown in Fig. 1. This visualization also provides insights into the scale of the generated libraries and the corresponding weight yields within a narrow energy range. Our strategy is to run any high-level analysis with the two productions and to calculate the desired final kinematic distribution by subtracting them. Furthermore, in addition to building a comprehensive statistical library of GiBUU events, we apply interaction vertex rotation techniques to avoid event duplication. Figure 2 shows the neutrino energy distribution for the two productions in the LArBox geometry. The ”negative weights” curve is significantly lower than the ”positive weights” at these energies.
Figure 1: Distribution of weights in the GiBUU library generated for this study for muon neutrino charged-current interaction for neutrino energy in $0.600-0.601$ GeV.
Figure 2: Neutrino energy distribution in the LArBox geometry for the ”negative” and ”positive” weights.
Our code is a customized version of the GENIE evtLib [10], which diverges from the original in several aspects. Notably, our customized evtLib can handle weighted events and crucially preserves the event identity, retaining the information related to the event generator reaction code. Subsequently, these events are incorporated into the LArSoft framework as an external library. The flux and detector geometry drivers developed within the GENIE framework are employed during this stage. The same methodology could be used for other cross-section models not yet available as event generators, such as some Beyond the Standard Model theory simulations. Our simulation study employed the described technique to simulate GiBUU events within the LArBox geometry, utilizing the BNB (Booster Neutrino Beam) neutrino beam [8] prediction.
Following this, we conducted several validation procedures to check the authenticity of the produced events by comparing them to the original GiBUU cross-section library, ensuring their conformity to the original GiBUU model. Figure 3 shows two of these studies looking at $\nu_{\mu}$ charged-current ($\nu_{\mu}$ CC) events with respect to the neutrino energy: the average hadron multiplicity and the ratio between events with a MEC-enhanced selection (one muon, two protons, and zero pions in the final state) and a QE-enhanced selection (one muon, one proton, and zero pions in the final state). Besides statistical fluctuations, especially at larger energies (where the BNB flux at LArBox is negligible) and minor differences due to the geometrical effects, the validation was successfully completed. The validation tests confirmed the GiBUU model’s transformation into a Monte Carlo event generator incorporating the appropriate flux and geometry settings.
Figure 3: Validation plots comparing the GiBUU cross-section library vs. GiBUU events generated in the LArBox geometry. The plot on the top shows the average hadron multiplicity per $\nu_{\mu}$ CC interactions with respect to the neutrino energy. The bottom plot calculates the ratio between a selection that enhances MEC (one muon, two protons, and zero pions) and a selection that enhances QE interactions (one muon, one proton, and zero pions).
IV Unaddressed Developments in Systematical Uncertainty Estimation for GiBUU
Like any model, the GiBUU framework is subject to systematic uncertainties that can impact the precision of its predictions. After converting GiBUU into a Monte Carlo, the next step is to assess these systematic uncertainties and develop the software framework necessary to propagate those uncertainties. This section aims to delve into the areas of systematic uncertainty estimation that need to be designed for GiBUU. The inherent uncertainties associated with nuclear physics inputs can introduce systematic biases in the simulation results. Understanding and quantifying these uncertainties is crucial for robust data analysis and extracting meaningful physics results.
We observe various sources contribute to systematic uncertainties in the GiBUU model:
Nuclear Effects and Medium Modifications.
One of the potentially primary sources of systematic uncertainty in the GiBUU model arises from nuclear effects and medium modifications. While the model incorporates nuclear structure information and accounts for FSI, there is ongoing research to improve our understanding of these phenomena:
  • Nuclear density and structure: GiBUU relies on nuclear density distributions and form factors to describe the presence of nuclei. Variations in these parameters can introduce uncertainties in the simulation results.
  • Medium Modifications: GiBUU considers the modification of hadron properties within the nuclear medium. This feature includes hadron masses and width changes due to interactions with surrounding nucleons. A deeper exploration of the dynamics of these medium modifications is necessary to refine uncertainty estimates.
Resonance Contributions.
GiBUU incorporates resonance dynamics, but estimations on the uncertainties associated must be addressed:
  • Resonance Widths: The precise determination of resonance widths and their energy dependence remains challenging. Variations in the assumed resonance widths can lead to uncertainties in resonance production and decay contributions.
  • Coupling Constants: The coupling constants governing resonance interactions are parameters with associated uncertainties.
Neutrino-Nucleon Cross Sections.
Accurate cross-sections are crucial for reducing systematic uncertainties, but challenges persist:
  • Energy Dependence: Cross sections depend on energy, and uncertainties in these dependencies can propagate into the predictions.
  • Nuclear Environment: The cross sections within the nuclear medium differ from those in free space due to nuclear correlations and uncertainties in these models, especially in neutrino scattering, where constraining these uncertainties using neutrino data is more challenging.
These studies’ comparisons between GiBUU simulations and experimental data are crucial in assessing the systematic uncertainties and validating the model. We can identify some of the associated uncertainties by comparing simulated event observables to data from electron and hadron scattering experiments. However, comparisons to neutrino scattering data are also critical.
As neutrino physics advances and experimental data becomes increasingly precise, the need to reduce systematic uncertainties in theoretical models like GiBUU becomes more pressing. Collaborative efforts should exist to refine these model aspects, ultimately enhancing our ability to extract meaningful physics results.
V Results and Discussion
We performed studies within this work using LArBox comparing GiBUU and GENIE predictions regarding the momentum (Fig. 4) and the cosine of the angle (Fig. 5) of outgoing muons in charged-current muon neutrino interactions. These figures highlight the nuanced differences between the models when generating events using the BNB neutrino flux at the LArBox detector location and its geometry.
We calculated the ratio of different neutrino interaction topologies, such as $\nu_{\mu}CC0P0\pi^{\pm 0}$ (no protons or pions are observed in the final state above an energy threshold, defined as 50MeV in kinetic energy for protons and pions), $\nu_{\mu}CC1P0\pi^{\pm 0}$ (one proton and no pions are observed in the final state above the energy threshold), $\nu_{\mu}CCNP0\pi^{\pm 0}$ (more than one proton and no pions are observed in the final state above the energy threshold), $\nu_{\mu}CC1\pi^{\pm}$ (one charged pion, no neutral pions, and any number of protons are observed in the final state above the energy threshold), $\nu_{\mu}CC1\pi^{0}$ (one neutral pion, no charged pions, and any number of protons are observed in the final state above the energy threshold), $\nu_{\mu}CCOther$ (the remaining charged current events that didn’t fall into the previous categories), with respect to the inclusive $\nu_{\mu}CC$ events. The particle energy thresholds chosen are realistic if considering LArTPC detectors and large enough to reduce the impact due to the particle identification efficiency at lower energies.
Figure 4 shows the ratio between a given neutrino-argon topology and the inclusive sample with respect to the muon momentum for GiBUU at the top and GENIE at the bottom. The most noticeable difference is that the contributions of each topology are clearly different in each model. In GiBUU, the dominant contributions are $\nu_{\mu}CC1P0\pi^{\pm 0}$ and $\nu_{\mu}CC0P0\pi^{\pm 0}$, while in GENIE the main contributions are due to $\nu_{\mu}CCNP0\pi^{\pm 0}$ and $\nu_{\mu}CCOther$.
By investigating using different particle energy thresholds and looking at different particle multiplicity, it is observed that GENIE produces higher proton multiplicity than GiBUU; the differences can be explained by the different neutrino interaction compositions per model, particularly for MEC and resonant productions, but as well the nuclear medium modeling.
Figure 5 shows similar ratios but with respect to the cosine of the muon angle with respect to the neutrino direction. We observe certain suppression of events in the most forward muons that affect mostly the $\nu_{\mu}CC1P0\pi^{\pm 0}$ topology in GiBUU and the $\nu_{\mu}CCNP0\pi^{\pm 0}$ in GENIE. These distributions are of special importance to understanding neutrino data, and the different trends of each Monte Carlo offer much information.
The different compositions by Monte Carlo tell us about the potential background predictions we can find in experimental data, which are very different depending on the model choice. These different composition predictions would lead to important corrections in the neutrino energy reconstruction for an oscillation experiment and could lead to different answers.
Figure 4: Ratios between each exclusive topology and the inclusive charged-current muon neutrino predictions regarding the muon momentum comparing GiBUU and GENIE in the LArBox geometry using the BNB flux.
Figure 5: Ratios between each exclusive topology and the inclusive charged-current muon neutrino predictions regarding the muon angle comparing GiBUU and GENIE in the LArBox geometry using the BNB flux.
In high-resolution detectors, such as LArTPC, the final particle content observed in the detector is critical for reconstructing neutrino energy. This approach, aimed at reducing model-dependent uncertainties compared to methods such as CCQE-like approaches, poses a significant challenge in identifying all hadrons emerging from the nucleus. A major hurdle is the quantification of invisible neutrons and the complex modeling of final proton content and their kinematics. Correcting for these observables, especially for nucleons, can introduce substantial uncertainties and biases in neutrino energy reconstruction. Accurate modeling of the number of nucleons knocked out during neutrino interactions is critical. Access to diverse modeling approaches is essential for addressing this challenge. In Figure 6, we present the average number of protons generated per neutrino interaction across varying neutrino energies. Solid lines represent predictions from the GiBUU model, while dashed lines correspond to GENIE predictions. This comparison underscores the nuanced differences between these models that will directly reflect when reconstructing neutrino energies at experiments such as SBN and DUNE.
Figure 6: Average number of protons simulated above 50 MeV per neutrino at each neutrino energy, per interaction type. Solid lines represent predictions from GiBUU; dashed lines represent predictions from GENIE.
The observed variations emphasize the significance of employing multiple models, such as GiBUU, to interpret neutrino scattering data accurately. These differences prompt further investigation into specific aspects of neutrino-nucleus interactions, refining our understanding and mitigating model-dependent uncertainties.
Validation against experimental data, including charged-current and neutral-current interactions, will demonstrate the model’s capacity to reproduce a broad range of observables faithfully. Our study reinforces the importance of model diversity in refining the interpretation of neutrino scattering data, which is essential for pushing the boundaries of our understanding in this intricate field.
VI Conclusion
In conclusion, our implementation of the GiBUU model as a Monte Carlo has yielded valuable insights into the complex dynamics of neutrino scattering processes. This Monte Carlo reproduces correctly the GiBUU model, and it is the first step in incorporating uncertainties to be propagated into experimental data analysis. Notably, a comparative study with GENIE reveals nuanced differences, emphasizing the necessity of employing multiple models to interpret neutrino scattering data comprehensively.
While acknowledging tensions in specific datasets, GiBUU’s overall robust performance establishes it as a reliable tool in the neutrino physics toolkit. The presented results underscore the importance of scrutinizing and refining model predictions, essential for advancing our comprehension of neutrino-nucleus interactions.
VII Acknowledgments
We thank Prof. Ulrich Mosel and Prof. Kai Gallmeister, from Justus-Liebig-Universitat Giessen, Germany, for their indispensable guidance. Their expertise significantly enhanced our study of neutrino interactions.